# Hype, Sustainability, and the Price of the Bigger-is-Better Paradigm in AI

### v1

@@ 2.3 Unrepresentative benchmarks show outsize benefits from scale
Furthermore, with the advent of generative models including LLMs, the ML community has faced new evaluation challenges given the openendedness of model outputs. It is no longer possible to simply evaluate prediction on labels in the test set when there is no canonical ‘correct answer’. This has resulted in a diversification of model evaluation practices, with the development of many different evaluation benchmarks intended to evaluative different aspects of generative model performance (see Koch et al., 2021; Chang et al., 2023). This proliferation of non-standard approaches further confuses the landscape of assessment. Furthermore, examining training datasets often reveals evidence of data contamination (Dodge et al., 2021), including the presence of benchmark datasets inside training corpora (Deng et al., 2023). This renders evaluations against these benchmarks questionable. This suggests that we should be very cautious regarding the validity of current evaluations of industrial models, given that the datasets used to train them are often not accessible to carry out the kind of public scrutiny necessary to identify such contamination or to understand the suitability of a given evaluation approach to the model at hand.
These issues have led researchers to propose various approaches for more rigorous model evaluation, including libraries (Gao et al., 2023; Von Werra et al., 2022) and open leaderboards (Beeching et al., 2023). However, the crux of the matter remains that ML benchmarks and evaluations remain unrepresentative proxies of downstream model performance, and are not a substitute for more holistic assessments. Importantly, such benchmarks cannot speak to whether a given model is suited to a given purpose and context. In addition, the “cost functions” that are used for many of these comparisons are not absolute and result in, e.g., the fixation on specific metrics such as accuracy or precision, while ignoring others, such as robustness or efficiency, which remain crucial in real-world applications (Rogers & Luccioni, 2024). Finally, some ad-hoc benchmarks claim to measure ill-defined performance properties that are in fact impossible to verify (e.g. claiming that certain LLMs exhibit artificial general intelligence (Bubeck et al., 2023)), which impacts the way the field as a whole tracks and perceives progress towards an illusory and unattainable goal.
@@ 3 Consequence 1: An unsustainable trajectory
As we have shown in the sections above, focusing all of our research efforts on the pursuit of bigger models narrows the field, and scale is not an efficient solution to every problem. There are many interesting problems in AI that do not require scale, and many different ways of measuring performance that do not only emphasize performance. The singular pursuit of scale also comes with worrying consequences beyond the ML research community, which we examine below.
@@
As we have seen from historical data, the compute required to create and deploy SOTA AI models grows faster than the cost of compute decreases. And yet, if we succeed at making these technologies generally useful, deployment at massive scale would be the logical next step, making these SOTA models accessible to increasing quantities of users. We may hope that efficiency improvements will come to the rescue, but the recent history of AI suggests the opposite. It is well-known phenomenon in economics that when the efficiency of a general-use technology increases, the falling costs lead to an increased demand, resulting in an overall increase in resource usage– this is referred to as “Jevons Paradox” (Jevons, 1865). Indeed, the AlexNet paper, which helped launch the current bigger-is-better era of AI research, was itself introducing a profound efficiency improvement, showing that a few GPUs could compete with tens of thousands of CPUs. One could assume that this would make AI training “democratically accessible”, but in fact the opposite happened, with subsequent generations of AI models using even larger quantities of GPUs and training for a longer amount of time – with the training of the most recent generations of LLMs being measured in millions of GPU hours. This counter-intuitive dynamic has been documented numerous times in energy, agriculture, transportation, and computing  (York & McGee, 2016), where observed trends have led to truisms such as Wirth’s law, which asserts that software demands outpace hardware improvements (Wirth, 1995).
@@ 3.1 The costs of scale-first AI don’t add up
In fact, in many organizations, compute costs are already a major roadblock to developing and deploying ML models (Paleyes et al., 2022). This is true even at large, tech-savvy companies such as Booking.com, who have found that gains in model performance cease to translate into gains in value to the company at a certain point, at which other factors such as efficiency or robustness become more important (Bernardi et al., 2019). Increase in model scale aggravates this problem – even after the costs of training are accounted for, deploying large models can lead to sizeable costs. In fact, the price of a single inference (i.e. model query) has been growing in the last two decades despite hardware efficiency improvements (Figure 3). This problem is made even worse as growing applications of such models in user-facing applications such as smartphones, smart devices connected to the Internet of Things (IOT) and Web applications require increased quantities of inferences by ML models.
@@
The prohibitive cost of model deployment at scale threatens the business model of AI. For instance, the compute required to run ChatGPT as a publicly accessible interface allegedly costs OpenAI \$ 700 000 per day (The Economic Times, 2023). According to Alphabet’s chairman, a search using Bard, which is powered by Google’s recent PaLM 2 model (Anil et al., 2023), is 10 times more expensive than a pre-Bard Google search (Dastin & Nellis, 2023). In a similar vein, Bornstein et al. (2023) reports that in generative AI gross margins are “more often as low as 50-60%, driven largely by the cost of model inference”. Beyond generative AI, self-driving cars provide another cautionary example. But here again the economics are challenging – ‘autonomous’ driving has turned out to rely heavily on the support of human workers and expensive sensors (Higgins, 2022). As of early 2024, the field of companies endeavoring to commercialize fully autonomous vehicles has shrunk considerably. Cruise suspended its efforts, leaving only Waymo remaining (Shepardson & Klayman, 2023).
@@ 3.2 Worrying environmental impacts
The emphasis on scale in AI also comes with consequences to the planet, since training and deploying AI models requires raw materials for manufacturing computing hardware, energy to power infrastructure, and water to cool ever-growing datacenters. The energy consumption and carbon footprint of AI models has grown with models such as LLMs emitting up to 550 tonnes of CO2 during their training process  (Strubell et al., 2019; Luccioni & Hernandez-Garcia, 2023; Luccioni et al., 2022) . But the most serious sustainability concerns relate to inference, given the speed at which AI models are being deployed in user-facing applications. It is hard to gather meaningful data regarding both the energy cost and carbon emissions of AI inference because this varies widely depending on the deployment choices made (e.g. type of GPU used, batch size, precision, etc.). However, high-level estimates from Meta attribute approximately one-third of their AI-related carbon footprint to model inference (Wu et al., 2021), whereas Google attributes 60% of its AI-related energy use to inference (Patterson et al., 2022). Comparing estimates of the energy required for inference for LLMs (Luccioni et al., 2024) to ChatGPT’s 10 million users per day (Oremus, 2023) reveals that within a few weeks of commercial use, the energy use of inference overtakes that of training. Given the many companies and investors working to make AI more ubiquitous, these numbers could go up by many orders of magnitude. In fact, we are already seeing chatbots applied in consumer electronics such as smart speakers and TVs, as well as appliances such as refrigerators and ovens. The increased use of AI in user-facing products is putting tech companies’ climate targets at risk, with both Microsoft and Google announcing in 2024 that they would miss the sustainability targets they set in previous years due to the energy demands of AI (Rathi & Bass, 2024; Metz, 2024).
@@
Comparing different model architectures also shows that “general-purpose” (multi-task or zero-shot) models have higher energy costs than specialized models that were trained or fine-tuned for a specific task (Luccioni et al., 2024, and Figure 4). This means that the current trend of using generative AI approaches for tasks such as Web search, which were previously carried out using “legacy” (information retrieval and embedding-based) systems, stands to increase the environmental costs of our day-to-day computer use significantly. Taking a step back, according to recent figures, global data centre electricity consumption represents 1-1.3% of global electricity demand and contributes 1% of energy-related greenhouse gas emissions (Hintemann & Hinterholzer, 2022; Copenhagen Centre on Energy Efficiency, 2020). It is hard to estimate what portion of this number is attributable to AI. However, a recent report from the International Energy Agency estimates that electricity consumption from data centres and AI is set to double in the next 2 years, with data centres’ total electricity consumption surpassing that of Japan (1 000 TWh) in 2026. (IEA, 2024).
@@ 4 Consequence 2: More data, more problems
As ML datasets grow in size (Figure 5), they bring a slew of issues ranging from documentation debts to a lack of auditability to biases – we discuss these below.
@@
In recent years, pretraining has become the dominant approach in both computer vision (Szegedy et al., 2015; Redmon et al., 2016) and natural language processing (Devlin et al., 2018; Liu et al., 2019). This approach requires access to large datasets, which have grown in size, from millions of images for datasets such as ImageNet (Deng et al., 2009) to gigabytes of textual documents for C4 (Raffel et al., 2020) and billions of image-text pairs in LAION-5B (Schuhmann et al., 2022). The premise of pretraining is that more data will improve model performance and ensure maximal coverage in model predictions, with the assumption that the more data used for pretraining, the more representative the model’s coverage will be of the world at large. However, numerous studies have shown that neither image nor text datasets are broadly representative, reflecting instead a select set of communities, regions and populations (Dodge et al., 2021; Rogers, 2021; Raji et al., 2021; Luccioni & Rolnick, 2023). In fact, recent research on the LAION datasets showed that as these datasets grow in size, the issues that plague them also multiply, with larger datasets contain disproportionally more problematic content than smaller datasets (Birhane et al., 2023). Thiel (2023) reported that the LAION datasets included child sexual abuse material, prompting the datasets to be taken down from several hosting platforms. But with dozens of image generation models trained on LAION variants, some already deployed in user-facing settings, it is hard to assess and mitigate the negative effects of this grim reality (Sambasivan et al., 2021).
While a growing wave of research has proposed a more ‘data-centric’ approach to data collection and curation (Zha et al., 2023; Mitchell et al., 2022), attempts to actually document the contents of ML datasets have been hampered by their sheer size, which requires compute and storage beyond the grasp of most members of the ML research community (Luccioni & Viviano, 2021), before the challenges of classifying and understanding the contents of a given dataset are even addressed. As the field works with ever larger datasets, we are also incurring more and more ‘documentation debt’ wherein training datasets are too large to document both during creation and post-hoc (Bender et al., 2021). In a nutshell, this means that we do not truly know what goes in to the models that we rely on to answer our questions and generate our images, apart from some high-level statistics. And this, in turn, hampers efforts to audit, evaluate, and understand these models.
@@
From a privacy perspective, the perceived need for ever-growing datasets implicitly incentivizes more pervasive surveillance, as companies gather our clicks, likes and searches to feed models that are then applied to sell us consumer products or impact the order the search results that we see. In fact, the majority of the revenue of the Big Tech companies leading the AI revolution comes from targeted advertising - representing 80% of Google’s and 90% of Meta’s annual revenue in 2022.
Also, while much of ML data gathering efforts have been operating under the assumption that copyright laws do not apply for data gathered from the Internet and used to train ML models (or, at the least, that training ML models constitutes ‘fair use’), recent months have seen a series of copyright lawsuits filed against many companies and organizations. This includes lawsuits from authors, artists and newspapers (Small, 2023; Schrader, 2023; Grynbaum & Mac, 2023). While is is too early to tell what the verdicts of these lawsuits will be, they will undoubtedly impact on the way in which ML datasets are created and leveraged. The European Union’s General Data Protection Regulation (GDPR) law sets some limitations in terms of data gathering and privacy in the EU, even if unevenly enforced. However, the United States and many other jurisdiction have yet to pass federal privacy laws, offering their citizens little to no privacy protection against predatory data gathering practices and surveillance, which the rush to scale AI incentivises.
@@ 5.2 Scale comes with a Concentration of power
For training and deploying large models, most AI practitioners operating outside of large cloud companies must rent access to compute from these companies. Since the cost of training ML systems has grown exponentially, this can entail costs for large scale models in the tens to hundreds of millions of dollars, which is beyond the budget of the vast majority of academic researchers (Figure 6). However, tech companies are more able to invest this money to fuel their AI research and competitiveness – Meta alone is estimated to have spent \$18 billion on GPUs in 2024 (Hays, 2024). While AI startups continue to raise significant capital, the majority of this capital (“up to 80-90% in early rounds”) is paid to cloud providers (Bornstein et al., 2023). Notably, the flow of capital here is often circular: three large cloud companies “contributed a full two-thirds of the \$27bn raised by fledgling AI companies in 2023” (Hammond, 2024).
These circular investment deals can blur the boundary between “investment” and “acquisition”. For instance, the deal between Microsoft and OpenAI (currently under investigation by the Federal Trade Commission, 2024), involved providing OpenAI access to Azure compute, in exchange for exclusive license to integrate OpenAI’s GPT models, and a promise of \$1 trillion in profit delivered to Microsoft prior to any revenue being directed to OpenAI’s social mission.
@@
AI dominance is one factor that has helped vault companies into shapers of global economic forecasts and markets. The S&P 500, used by US investors as an index of market performance, is significantly shaped by the fates of large US tech companies (Rennison & Murray, 2023). This indicates that the interests of large AI companies, and particularly their emphasis on large-scale AI, have implications well beyond the tech industry, since a downturn in the AI market would impact global financial markets beyond. This creates market incentives, divorced from scientific imperatives, for perpetuating the bigger-is-better paradigm on which these companies are betting.
The expense of creating large-scale AI models also makes the need to commercialize these models more pressing – even if inference can be costly (Vanian & Leswing, 2023). Here, established cloud and platform companies also have an advantage in the form of access to large and existing markets. Cloud infrastructure offerings are a readily available means to commercialize AI models, allowing startups and enterprises to rent access to, e.g., generative AI APIs as part of cloud contracts. The ecosystem of ‘GPT wrapper’ companies licensing access to the model via Microsoft’s Azure cloud is an obvious example. McKinsey (2022), looking at the 2022 landscape of AI, noted that “commercialization is likely tied to hosting. Demand for proprietary APIs (e.g. from OpenAI) is growing rapidly.” Notable here is that the business model for commercializing generative AI in particular is still being worked out (Bornstein et al., 2023).
@@
The 1990s were marked by the explosion of personal computing, and the race to commercialize networked computation; this innovation reshaped the economy and the labor market (Caballero, 2010). From there, the actors involved became concentrated into a small number of firms, becoming enablers and gatekeepers of personal-computing software. Resulting market capture created a rent, to the expense of the rest of the economy, including IT consumers (Aghion et al., 2023). This, in turn, led to a fall of innovation and growth because of lack of incentives for leaders and lack of market access for competitors (Aghion et al., 2023); desktop software has been moving slowly.
@@
The growth and profit imperative of corporations dictates finding sources of revenue that can cover the costs of AI systems of ever-growing size. These often puts them at odds with responsible and ethical AI use, which puts the emphasis on more consensual and incremental progress. We see this dynamic at play in OpenAI’s shifting stance and policies as they have become more closely tied to Microsoft, a choice they made admittedly to access scarce computational resources. This is illustrated by the 2024 change to their acceptable use policy, which significantly softened the proscription against using their products for “military and warfare”; this change unlocks new revenue streams (Stone & MBergen, 2024), and connects OpenAI’s large scale AI to Microsoft’s existing US military partnerships (Xiang, 2023). The public did not have a say in this determination, nor, we assume, did the global stakeholders whose interests may not be safeguarded by the militaries that OpenAI chooses to provide services to. And yet, the concentrated private industry power over AI creates a small, and financially incentivized segment of AI decision makers. We should consider how such concentrated power with agency over centralized AI could shape society under more authoritarian conditions.
There are many other examples of financial incentives shaping the use of automated decision systems in ways that are not socially beneficial. Ross & Herman (2023) showed United Health used a diagnostic algorithm to deny patients care, an application now at the center of a class action lawsuit filed by patients and their families. The current bigger-is-better AI paradigm is also exacerbating geopolitical concentration and tensions. Beyond the power of large corporations, government imperatives are also in play, particularly given that the large AI industry players are primarily located in the US. These tensions are apparent in the efforts on the part of the US government to limit the supply of computing resources such as GPUs to China in the global race for AI dominance (Alper et al., 2023). While a discussion of whether such limits are warranted is outside the scope of this paper, this dynamic highlights the role of AI as a source of geopolitical power, one that given the current concentration of AI resources threatens to undermine technological sovereignty globally.
@@
Deploying AIs is society may come with a variety of risks. Some risks are already visible in existing systems such as the problems of fairness and replicating biases which lead to underdiagnosis in historically under-served populations (Seyyed-Kalantari et al., 2021). Other risks are more speculative given today’s systems but are to be considered in future scenario where AIs become much more powerful (Critch & Russell, 2023). A concentration of power fostered by very large-scale AI systems that can be operated and control by only very few actors makes these multiple risks more problematic, as it challenges checks and balance mechanisms.

### v2

@@ 2.3 Unrepresentative benchmarks show outsize benefits from scale
Furthermore, with the advent of generative models including LLMs, the ML community has faced new evaluation challenges given the openendedness of model outputs. It is no longer possible to simply evaluate prediction on labels in the test set when there is no canonical ‘correct answer’. This has resulted in a diversification of model evaluation practices, with the development of many different benchmarks intended to evaluative different aspects of generative model performance (Koch et al., 2021; Chang et al., 2023) – this proliferation of non-standard approaches further confuses the landscape of assessment. To make matters worse, examining training datasets often reveals evidence of data contamination (Dodge et al., 2021), including the presence of benchmark datasets inside training corpora (Deng et al., 2023). This renders evaluations against these benchmarks questionable. All these shortcomings in benchmarks suggest that we should be very cautious regarding the validity of current evaluations of industrial models, given that the datasets used to train them are often not accessible to carry out the kind of public scrutiny necessary to identify such contamination or to understand the suitability of a given evaluation approach to the model at hand.
These issues have led to research on more rigorous model evaluation, including libraries (Gao et al., 2023; Von Werra et al., 2022) and open leaderboards (Beeching et al., 2023). However, the crux of the matter remains that ML benchmarks and evaluations remain unrepresentative proxies of downstream model performance, and are not a substitute for more holistic assessments. Importantly, such benchmarks cannot speak to whether a given model is suited to a given purpose and context. In addition, the “cost functions” used for many of these comparisons are not absolute and result in, e.g., the fixation on specific metrics such as accuracy or precision, while ignoring others, such as robustness or efficiency, which remain crucial in real-world applications (Rogers and Luccioni, 2024). Finally, some ad-hoc benchmarks claim to measure ill-defined performance properties that are in fact impossible to verify (e.g. claiming that certain LLMs exhibit artificial general intelligence (Bubeck et al., 2023)). Considering a set of model capabilities much more narrow and better defined that general intelligence, model reasoning, thorough inspection of model capabilities has shown that benchmarks cannot be trusted as they are full of shortcuts exploited by models (Mirzadeh et al., 2024). Such frail measures of high-level model capabilities impact the way the field as a whole tracks and perceives progress towards an ill-defined, illusory, and unattainable goal of general intelligence.
@@ 3 Consequence 1: An unsustainable trajectory
As we have shown in the sections above, focusing all of our research efforts on the pursuit of bigger models narrows the field, and scale is not an efficient solution to every problem. There are many interesting problems in AI that do not require scale. The singular pursuit of scale also comes with worrying consequences, which we examine below.
@@
As we have seen from historical data, the compute required to create and deploy SOTA AI models grows faster than the cost of compute decreases. And yet, if we succeed at making these technologies generally useful, deployment at massive scale would be the logical next step, making these SOTA models accessible to increasing quantities of users. We may hope that efficiency improvements will come to the rescue, but the recent history of AI suggests the opposite. It is well-known phenomenon in economics that when the efficiency of a general-use technology increases, the falling costs lead to an increased demand, resulting in an overall increase in resource usage. This is referred to as “Jevons Paradox” (Jevons, 1865). Indeed, the AlexNet paper, which helped launch the current bigger-is-better era of AI research, was itself introducing a profound efficiency improvement, showing that a few GPUs could compete with tens of thousands of CPUs. One could assume that this would make AI training “democratically accessible”, but in fact the opposite happened, with subsequent generations of AI models using even larger quantities of GPUs and training for a longer amount of time – with the training of the most recent generations of LLMs being measured in millions of GPU hours. This counter-intuitive dynamic has been documented numerous times in energy, agriculture, transportation, and computing  (York and McGee, 2016), where observed trends have led to truisms such as Wirth’s law, which asserts that software demands outpace hardware improvements (Wirth, 1995).
@@ 3.1 The costs of scale-first AI don’t add up
In fact, in many organizations, compute costs are already a major roadblock to developing and deploying ML models (Paleyes et al., 2022). This is true even at large, tech-savvy companies such as Booking.com, who have found that gains in model performance cease to translate into gains in value to the company at a certain point, at which other factors such as efficiency or robustness become more important (Bernardi et al., 2019). Increase in model scale aggravates this problem: even after the costs of training are accounted for, large models come with large deployment costs. Indeed, the price of a single inference (i.e. model query) has been growing in the last two decades despite hardware efficiency improvements (Figure 3).
@@
Prohibitive costs of model deployment at scale threatens the business model of AI. Due to these costs, major investors are questioning the financial health of the current AI massive growth (Cahn, 2024; Acemoglu et al., 2024). For instance, the compute required to run ChatGPT as a publicly accessible interface cost OpenAI \$ 700 000 per day in 2023 (The Economic Times, 2023), and OpenAI’s 2024 estimated balance is a \$ 5 B loss (Isaac and Griffith, 2024). According to Alphabet’s chairman, a search using Bard, which is powered by Google’s recent PaLM 2 model (Anil et al., 2023), is 10 times more expensive than a pre-Bard Google search (Dastin and Nellis, 2023). In a similar vein, Bornstein et al. (2023) reports that in generative AI gross margins are “more often as low as 50-60%, driven largely by the cost of model inference”. The problem of costs is increasingly pressing and many recent announcements around industry R&D have focused on the cost or energy efficiency of new developments: OpenAI announced their recent GPT-4o mini as “advancing cost-efficient intelligence” (OpenAI, 2024). DeepSeek-v3 (Liu et al., 2024) is a very large (671B parameters) and capable language model which positions itself putting forward much decreased training and inference costs. The modernBERT paper (Warner et al., 2024) tackles language processing while specifically avoiding generative modeling to decrease costs. Microsoft CEO Satya Nadella even argued in late 2024 that progress in generative AI should be measured in “token per dollar per watt”, as opposed to metrics focusing only on accuracy- or cost (Grallet, 2024). As models are improving, the costs of one query are going down for a given model size. However, will this be enough? “GenAI” product development is calling not only for increased model scale –eg to alleviate hallucinations–, but also increasing quantities of inferences by these models, adding them in user-facing applications on the web, but also in smartphones, smart connected devices, cars, etc.
Beyond generative AI, self-driving cars provide cautionary example. Here again the economics are challenging – ‘autonomous’ driving has turned out to rely heavily on the support of human workers and expensive sensors (Higgins, 2022). As of 2024, the field of companies endeavoring to commercialize fully autonomous vehicles has shrunk considerably. Cruise suspended its efforts, leaving only Waymo remaining (Shepardson and Klayman, 2023).
@@ 3.2 Worrying environmental impacts
The emphasis on scale in AI also comes with consequences to the planet, since training and deploying AI models requires raw materials for manufacturing computing hardware, energy to power infrastructure, and water to cool ever-growing datacenters. The energy consumption and carbon footprint of AI models has grown with models such as LLMs emitting up to 550 tonnes of CO2 during their training process  (Strubell et al., 2019; Luccioni and Hernandez-Garcia, 2023; Luccioni et al., 2022) . But the most serious sustainability concerns relate to inference, given the speed at which AI models are being deployed in user-facing applications. It is hard to gather meaningful data regarding both the energy cost and carbon emissions of AI inference because this varies widely depending on the deployment choices made (e.g. type of GPU used, batch size, precision, etc.). However, high-level estimates from Meta attribute approximately one-third of their AI-related carbon footprint to model inference (Wu et al., 2021), whereas Google attributes 60% of its AI-related energy use to inference (Patterson et al., 2022). Comparing estimates of the energy required for inference for LLMs (Luccioni et al., 2024) to ChatGPT’s 10 million users per day (Oremus, 2023) reveals that within a few weeks of commercial use, the energy use of inference overtakes that of training. Given the efforts to make AI more ubiquitous, these numbers could go up by many orders of magnitude. In fact, we are already seeing chatbots applied in consumer electronics such as smart speakers and TVs, as well as appliances such as refrigerators and ovens. The increased use of AI in user-facing products is putting tech companies’ climate targets at risk, with both Microsoft and Google announcing in 2024 that they would miss the sustainability targets they set in previous years due to the energy demands of AI (Rathi and Bass, 2024; Metz, 2024).
@@
Comparing different model architectures also shows that “general-purpose” (multi-task or zero-shot) models have higher energy costs than specialized models that were trained or fine-tuned for a specific task (Luccioni et al., 2024, and Figure 4). This means that the current trend of using generative AI approaches for tasks such as Web search, which were previously carried out using “legacy” (information retrieval and embedding-based) systems, stands to increase the environmental costs of our day-to-day computer use significantly. The impact of a models’ inference may seem a trifle compared to say a plane ride, however here again the challenge comes from the huge number of times they can be used. Data centers are already putting noticeable pressure on electricity networks of modern countries such as the US (Nicoletti et al., 2024), leading tech companies to fund nuclear power plants for their exclusive usage (Mandler, 2024). Taking a step back, according to recent figures, global data centre electricity consumption represents 1-1.3% of global electricity demand and contributes 1% of energy-related greenhouse gas emissions (Hintemann and Hinterholzer, 2022; Copenhagen Centre on Energy Efficiency, 2020). It is hard to estimate what portion of this number is attributable to AI. However, a recent report from the International Energy Agency estimates that electricity consumption from data centres and AI is set to double in 2 years, surpassing that of Japan (1 000 TWh) in 2026. (IEA, 2024).
@@ 4 Consequence 2: More data, more problems
As ML datasets grow in size (Figure 5), they bring a slew of issues ranging from documentation debts to a lack of auditability to biases. We discuss these below.
@@
In recent years, pretraining has become the dominant approach in both computer vision (Szegedy et al., 2015; Redmon et al., 2016) and natural language processing (Devlin et al., 2018; Liu et al., 2019). This approach requires access to large datasets, which have grown in size, from millions of images for datasets such as ImageNet (Deng et al., 2009) to gigabytes of textual documents for C4 (Raffel et al., 2020) and billions of image-text pairs in LAION-5B (Schuhmann et al., 2022). The premise of pretraining is that more data will improve model performance and ensure maximal coverage in model predictions, with the assumption that the more data used for pretraining, the more representative the model’s coverage will be of the world at large. However, numerous studies have shown that neither image nor text datasets are broadly representative, reflecting instead a select set of communities, regions and populations (Dodge et al., 2021; Rogers, 2021; Raji et al., 2021; Luccioni and Rolnick, 2023). In fact, recent research on the LAION datasets showed that as these datasets grow in size, the issues that plague them also multiply, with larger datasets contain disproportionally more problematic content than smaller datasets (Birhane et al., 2023). Thiel (2023) reported that the LAION datasets included child sexual abuse material, prompting the datasets to be taken down from several hosting platforms. But with dozens of image generation models trained on LAION variants, some already deployed in user-facing settings, it is hard to assess and mitigate the negative effects of this grim reality (Sambasivan et al., 2021).
While a growing wave of research has proposed a more ‘data-centric’ approach to data collection and curation (Zha et al., 2023; Mitchell et al., 2022), attempts to actually document the contents of ML datasets have been hampered by their sheer size, which requires compute and storage beyond the grasp of most members of the ML research community (Luccioni and Viviano, 2021), before the challenges of classifying and understanding the contents of a given dataset are even addressed. As the field works with ever larger datasets, we are also incurring more and more ‘documentation debt’ wherein training datasets are too large to document both during creation and post-hoc (Bender et al., 2021). In a nutshell, this means that we do not truly know what goes in to the models that we rely on to answer our questions and generate our images, apart from some high-level statistics. And this, in turn, hampers efforts to audit, evaluate, and understand these models.
@@
From a privacy perspective, the perceived need for ever-growing datasets implicitly incentivizes more pervasive surveillance, as companies gather our clicks, likes and searches to feed models that are then applied to sell us consumer products or impact the order the search results that we see. In fact, the majority of the revenue of the Big Tech companies leading the AI revolution comes from targeted advertising - representing 80% of Google’s (Alphabet, 2023) and 90% of Meta’s (Meta, 2023a, b) annual revenue in 2022.
Also, while much of ML data gathering efforts have been operating under the assumption that copyright laws do not apply for data gathered from the Internet and used to train ML models (or, at the least, that training ML models constitutes ‘fair use’), last year has seen a series of copyright lawsuits filed against many companies and organizations. This includes lawsuits from authors, artists and newspapers (Small, 2023; Schrader, 2023; Grynbaum and Mac, 2023). While is is too early to tell what the verdicts of these lawsuits will be, they will undoubtedly impact on the way in which ML datasets are created and leveraged. The European Union’s General Data Protection Regulation (GDPR) law sets some limitations in terms of data gathering and privacy in the EU, even if unevenly enforced. However, the United States and many other jurisdiction have yet to pass federal privacy laws, offering their citizens little to no privacy protection against predatory data gathering practices and surveillance, which the rush to scale AI incentivises.
@@ 5.2 Scale comes with a Concentration of power
For training and deploying large models, most AI practitioners operating outside of large cloud companies must rent access to compute from these companies. Since the cost of training ML systems has grown exponentially, this can entail costs for large scale models in the tens to hundreds of millions of dollars, which is beyond the budget of the vast majority of academic researchers (Figure 6). However, tech companies are more able to invest this money to fuel their AI research and competitiveness – Meta alone is estimated to have spent \$18 billion on GPUs in 2024 (Hays, 2024), while OpenAI announced early 2025 StarGate, a joint venture with SoftBank, Oracle, and MGX, planning more than \$ 100 billion AI-infrastructure investments (OpenAI, 2025), bigger than the GDP of most countries in the world. While AI startups continue to raise significant capital, the majority of this capital (“up to 80-90% in early rounds”) is paid to cloud providers (Bornstein et al., 2023). Notably, the flow of capital here is often circular: three large cloud companies “contributed a full two-thirds of the \$27bn raised by fledgling AI companies in 2023” (Hammond, 2024).
These circular investment deals can blur the boundary between “investment” and “acquisition”. For instance, the deal between Microsoft and OpenAI (currently under investigation by the FTC, Federal Trade Commission, 2024), involved providing OpenAI access to Azure compute, in exchange for exclusive license to integrate OpenAI’s GPT models, and a promise of \$1 trillion in profit delivered to Microsoft prior to any revenue being directed to OpenAI’s social mission.
@@
AI dominance is one factor that has helped vault companies into shapers of global economic forecasts and markets. The S&P 500, used by US investors as an index of market performance, is significantly shaped by the fates of large US tech companies (Rennison and Murray, 2023). This indicates that the interests of large AI companies, and particularly their emphasis on large-scale AI, have implications well beyond the tech industry, since a downturn in the AI market would impact global financial markets beyond. This creates market incentives, divorced from scientific imperatives, for perpetuating the bigger-is-better paradigm on which these companies are betting.
The expense of creating large-scale AI models also makes the need to commercialize these models more pressing – even if inference can be costly (Vanian and Leswing, 2023). Here, established cloud and platform companies also have an advantage in the form of access to large and existing markets. Cloud infrastructure offerings are a readily available means to commercialize AI models, allowing startups and enterprises to rent access to, e.g., generative AI APIs as part of cloud contracts. The ecosystem of ‘GPT wrapper’ companies licensing access to the model via Microsoft’s Azure cloud is an obvious example. McKinsey (2022), looking at the 2022 landscape of AI, noted that “commercialization is likely tied to hosting. Demand for proprietary APIs (e.g. from OpenAI) is growing rapidly.” Notable here is that the business model for commercializing generative AI in particular is still being worked out (Bornstein et al., 2023).
@@
When a few actors control an economy, it stifles innovation, as illustrated by recent history of computing. The 1990s were marked by the explosion of personal computing, and the race to commercialize networked computation; this innovation reshaped the economy and the labor market (Caballero, 2010). From there, the actors involved became concentrated into a small number of firms, becoming enablers and gatekeepers of personal-computing software. Resulting market capture created a rent, to the expense of the rest of the economy, including IT consumers (Aghion et al., 2023). This, in turn, led to a fall of innovation and growth because of lack of incentives for leaders and lack of market access for competitors (Aghion et al., 2023); desktop software has been moving slowly.
@@
The growth and profit imperative of corporations dictates finding sources of revenue that can cover the costs of AI systems of ever-growing size. These often puts them at odds with responsible and ethical AI use, which puts the emphasis on more consensual and incremental progress. We see this dynamic at play in OpenAI’s shifting stance and policies as they have become more closely tied to Microsoft, a choice they made admittedly to access scarce computational resources. This is illustrated by the 2024 change to their acceptable use policy, which significantly softened the proscription against using their products for “military and warfare”; this change unlocked new revenue streams (Stone and MBergen, 2024), connected OpenAI’s large scale AI to Microsoft’s existing US military partnerships (Xiang, 2023), and later pushed AI to direct battlefield application via a partnership with the defense-tech company Anduril (O’Donnell, 2024). The public did not have a say in this determination, nor, we assume, did the global stakeholders whose interests may not be safeguarded by the militaries that OpenAI chooses to provide services to. And yet, the concentrated private industry power over AI creates a small, and financially incentivized segment of AI decision makers. We should consider how such concentrated power with agency over centralized AI could shape society under more authoritarian conditions.
There are many other examples of financial incentives shaping the use of automated decision systems in ways that are not socially beneficial. Ross and Herman (2023) showed United Health used a diagnostic algorithm to deny patients care, an application now at the center of a class action lawsuit filed by patients and their families (US district court, district of Minnesota, 2023). The bigger-is-better AI paradigm is also exacerbating geopolitical concentration and tensions. Beyond the power of large corporations, government imperatives are also in play, particularly given that the large AI industry players are primarily located in the US. These tensions are apparent in the efforts on the part of the US government to limit the supply of computing resources such as GPUs to China in the global race for AI dominance (Alper et al., 2023). While a discussion of whether such limits are warranted is outside the scope of this paper, this dynamic highlights the role of AI as a source of geopolitical power, one that given the current concentration of AI resources threatens to undermine technological sovereignty globally.
@@
Deploying AIs is society may come with a variety of risks (Bengio et al., 2024). Some risks are already visible in existing systems such as the problems of fairness and replicating biases which lead to underdiagnosis in historically under-served populations (Seyyed-Kalantari et al., 2021). Other risks are more speculative given today’s systems but are to be considered in future scenario where AIs become much more powerful (Critch and Russell, 2023). A concentration of power fostered by very large-scale AI systems that can be operated and control by only very few actors makes these multiple risks more problematic, as it challenges checks and balance mechanisms.